the ph of a solution 11.5 calculate the hydrogen ion concentration of solution?

1
Expert's answer
2022-01-31T17:13:03-0500

1. pH= -lg[H+]

[H+]=10^-pH=10-11.5 3.16 × 10-12

2.pOH=14 - pH=14-11.5= 2.5

pOH=-lg[OH-];

[OH-]=10-2.5 3.16×10-3

Answer: 1. 3.16 × 10-12 M

2.3.16×10-3 M